# Observational constraints on inflationary models with non-minimally derivative coupling by ACT

### v1

@@ 1 Introduction
The results suggest that the inflationary models with higher values of $n_{s}$ and lower values of $r$ are favored. In particular, the ACT data appear to rule out monomial inflation models with convex potentials within the framework of Einstein’s general relativity (GR) [2]. These ACT findings have revitalized interest in alternative scenarios, including models predicting larger values of $n_{s}$ and those with nonminimal coupling [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
A widely studied class of inflationary models with nonminimal coupling involves a scalar field coupled to the Ricci scalar via a term of the form $f(\phi)R$, which is a specific case of the more general scalar-tensor theory $F(\phi,R)$ [41, 42, 43]. This type of coupling can be removed through a conformal transformation, reducing the theory to GR with a redefined scalar sector. However, if the kinetic term of the scalar field is non-minimally coupled to the curvature tensor, a conformal transformation cannot recast the model into a scalar-tensor theory [44]. Generally, non-minimally derivative couplings introduce higher than second-order derivatives in the field equations and extra degrees of freedom, potentially leading to the Boulware-Deser ghost [45]. A particular choice, where the derivative coupling takes the form $G^{\mu\nu}\phi_{,\mu}\phi_{,\nu}$ which is a special case of the most general Horndeski theory [46], avoids this issue because the resulting field equations in Horndeski theory contain no higher than second derivatives [46, 47]. Moreover, the gravitationally enhanced friction associated with this coupling slows down the evolution of the scalar field [48, 49, 50]. In this paper, we show that various inflationary models with such non-minimally derivative coupling are well consistent with the latest ACT data.
@@ 2 The model with non-minimally derivative coupling
The action for the scalar field with the kinetic term non-minimally coupled to Einstein tensor is
@@
where $M$ is the coupling constant with the dimension of mass and we take the Planck mass $M_{pl}^{2}=(8\pi G)^{-1}=1$. From the action (2.1), we get the Friedmann equation and the equation of motion for the scalar field $\phi$
@@
where $V_{\phi}=dV/d\phi$. To quantify those slow-roll conditions (2.4) in the high friction limit $H^{2}\gg M^{2}$, we introduce the following slow-roll parameters
@@
where $c_{s}$ and $c_{t}$ are the effective sound speed, $\nu\approx 3/2+4\epsilon-\eta$ and $\mu=3/2+\epsilon$. The amplitude of scalar power spectrum, the scalar spectral tilt and the tensor-to-scalar ratio are
@@
When $H^{2}\ll M^{2}$, we recover the results in GR.
@@
[…]
where the subscript “$e$” means the value of inflaton field at the end of inflation and the subscript “*” means the value at the horizon crossing. The value of $N$ is around $\sim 50-60$.
@@ 3.1 Power law potential
Note that $M^{2}\ll V_{0}$ in the high friction limit, without loss of generality, we take $M^{2}/V_{0}=10^{-4}$ in this paper. The end of inflation is determined by ${\rm max}\{\epsilon(\phi_{e}),|\eta(\phi_{e})|\}=1$. For $20<n<$ 2, inflation ends when $\epsilon(\phi_{e})=1$, so we have
@@
For $n\geq 2$, inflation ends when $\eta(\phi_{e})=1$, so
@@
where $\tilde{N}=N+n/2(n+2)$ for $20<n<$ 2 and $\tilde{N}=N+(n-1)/(n+2)$ for $n\geq 2$. Substituting Eqs. (3.3) and (3.5) into Eqs. (2.12) and (2.13), we obtain the scalar spectral tilt and the tensor-to-scalar ratio,
@@ 3.2 Hilltop model
Since $\phi<\mu$, so $\epsilon(\phi)$ is smaller than $\eta(\phi)$ by a factor $(\phi/\mu)^{p}$ when $p\geq 3$. In this model, we take $p=4$ as a representative case. By using Eqs. (2.12), (2.13) and (2.14), we numerically calculate the observables $n_{s}$ and $r$ with $\mu$ varying and the results are shown in Fig. 1. The solid blue lines represent the results for hilltop model with $p=4$ in the high friction limit, where the left and the right lines correspond to $N=50$ and $N=60$, respectively. In Fig. 1, we also show the results in GR with dashed blue lines. As seen from Fig. 1, the non-minimally derivative coupling increases $n_{s}$ and suppresses $r$, aligning the model with current observational constraints. For $N=50$ the model can be consistent with the P-ACT-LB-BK18 data only at the $2\sigma$ confidence level, while satisfying current observations at the $1\sigma$ confidence level for $N=60$. For $N=50$, we find that $2.40.18<\mu<$ 2.4 at the $2\sigma$ confidence level. For $N=60$, we find that $0.690.165<\mu<$ 0.69 at the $1\sigma$ confidence level and $\mu>0.123$ at the $2\sigma$ confidence level.
@@
As discussed above, $\epsilon(\phi)$ is smaller than $\eta(\phi)$ by a factor $(\phi/\mu)^{p}$ when $p\geq 3$, so inflation ends when $|\eta(\phi_{e})|=1$. In the limit of $\mu\ll 1$, we get small field inflation, $\phi\ll\mu$,
@@
$\begin{split}\epsilon(\phi)&=\frac{p^{2}(\phi/\mu)^{2p-2}}{2\mu^{2}\left[1-(\phi/\mu)^{p}\right]^{2}},\\ \eta(\phi)&=-\frac{p(p-1)(\phi/\mu)^{p-2}}{\mu^{2}\left[1-(\phi/\mu)^{p}\right]}=-\frac{2(p-1)}{p}\frac{1-(\phi/\mu)^{p}}{(\phi/\mu)^{p}}\epsilon(\phi).\end{split}$ (3.16)
When $p\geq 3$, the end of inflation is determined by $|\eta(\phi_{e})|=1$ The number of e-folds before the end of inflation is $N=g(\phi_{e}/\mu)-g(\phi_{*}/\mu)$, where
@@ 3.3 Polynomial $\alpha$-attractors
For the polynomial $\alpha$-attractors, the potential is
$V(\phi)=V_{0}\left(1-\frac{\mu^{p}}{\phi^{p}}\right),$ (3.18)
and the slow-roll parameters in the high friction limit are
@@
Without loss of generality, we take $p=3$. By using Eqs. (2.12), (2.13) and (2.14), we numerically calculate the observables $n_{s}$ and $r$ with $\mu$ varying, and the results are shown in Fig. 1 with solid magenta lines. With the help of non-minimally derivative coupling, the polynomial $\alpha$-attractors are consistent with the P-ACT-LB-BK18 data for both $N=50$ and $N=60$. For $N=50$, the $1\sigma$ and $2\sigma$ constraints on $\mu$ by the P-ACT-LB-BK18 data are $0.510.032<\mu<$ 0.51 and $3.560<\mu<$ 3.56, respectively; while for $N=60$, the $1\sigma$ constraint is $0.570<\mu<$ 0.57 and at the $2\sigma$ confidence level, $\mu$ can be arbitrarily large.
In the limit of $\mu\gg 1$, the same attractors (3.7) and (3.8) with $n=1$ are reached because the potential behaves as the linear potential when $\phi\sim\mu$. On the other hand, when $\mu\ll 1$, we have small field inflation. From $|\eta(\phi_{e})|=1$, we get
@@
$\begin{split}\epsilon(\phi)&=\frac{p^{2}(\mu/\phi)^{2p+2}}{2\mu^{2}\left[1-(\mu/\phi)^{p}\right]^{2}},\\ \eta(\phi)&=-\frac{p(p+1)(\mu/\phi)^{p+2}}{\mu^{2}\left[1-(\mu/\phi)^{p}\right]}=-\frac{2(p+1)}{p}\frac{1-(\mu/\phi)^{p}}{(\mu/\phi)^{p}}\epsilon(\phi).\end{split}$ (3.25)
@@ 3.4 Exponential $\alpha$-attractors
In this section, without loss of generality we consider $a=1$ for simplicity. The slow-roll parameters are
@@
By using Eqs. (2.12), (2.13) and (2.14), we numerically calculate $n_{s}$ and $r$ with varying $\alpha$ and the results are shown in Fig. 1. It is evident that the model is consistent with the P-ACT-LB-BK18 data at the $2\sigma$ confidence level with the help of non-minimally derivative coupling. For $N=50$, the $1\sigma$ and $2\sigma$ constraints on $\alpha$ are $0.0111.2\times 10^{-3}<\alpha<$ 0.011 and $0.4682.5\times 10^{-4}<\alpha<$ 0.468, respectively. For $N=60$, the $1\sigma$ and $2\sigma$ constraints are $0.0221.38\times 10^{-4}<\alpha<$ 0.022 and $\alpha>0$, respectively.
@@
In the small $\alpha$ limit, $\alpha\ll 1$, the end of inflation is determined by $|\eta(\phi_{e})|=1$, we have
@@
$\begin{split}\epsilon(\phi)&=\frac{1}{3\alpha\left[\exp\left(\sqrt{\frac{2}{3\alpha}}\phi\right)-1\right]^{2}},\\ \eta(\phi)&=\frac{2}{3\alpha\left[1-\exp\left(\sqrt{\frac{2}{3\alpha}}\phi\right)\right]}.\end{split}$ (3.35)
@@ 4 Conclusions
For the hilltop model, the scale $\mu$ can be smaller than the Planck energy, allowing for small-field inflation. In the case $p=4$, the scenario with $N=50$ lies within the $2\sigma$ $n_{s}-r$ contour from the P-ACT-LB-BK18 data, while the case with $N=60$ falls within both the $1\sigma$ and $2\sigma$ $n_{s}-r$ contours. For $N=50$, we constrain the scale as $2.40.18<\mu<$ 2.4 at the $2\sigma$ confidence level. For $N=60$, the bounds are $0.690.165<\mu<$ 0.69 at $1\sigma$, with a lower limit $\mu>0.123$ at $2\sigma$. Additionally, we derive analytical expressions for attractor behaviors in the limits of large and small $\mu$, for both the non-minimal derivative coupling scenario and GR.
For the polynomial $\alpha$-attractors with $p=3$, both the $N=50$ and $N=60$ cases are consistent with the P-ACT-LB-BK18 data. For $N=50$, the $1\sigma$ and $2\sigma$ constraints on $\mu$ are $0.510.032<\mu<$ 0.51 and $3.560<\mu<$ 3.56, respectively. For $N=60$, the $1\sigma$ constraint is $0.570<\mu<$ 0.57, while at the $2\sigma$ level, $\mu$ is unconstrained and can be arbitrarily large. The numerical results confirm the analytical attractor expressions for both the non-minimally derivative coupling scenario and GR in the limits of large and small $\mu$.

### v2

@@ 1 Introduction
The results suggest that the inflationary models with higher values of $n_{s}$ and lower values of $r$ are favored. In particular, the ACT data appear to rule out monomial inflation models with convex potentials within the framework of Einstein’s general relativity (GR) [2]. These ACT findings have revitalized interest in alternative scenarios, including models predicting larger values of $n_{s}$ and those with nonminimal coupling [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
A widely studied class of inflationary models with nonminimal coupling involves a scalar field coupled to the Ricci scalar via a term of the form $f(\phi)R$, which is a specific case of the more general scalar-tensor theory $F(\phi,R)$ [45, 46, 47]. This type of coupling can be removed through a conformal transformation, reducing the theory to GR with a redefined scalar sector. However, if the kinetic term of the scalar field is non-minimally coupled to the curvature tensor, a conformal transformation cannot recast the model into a scalar-tensor theory [48]. Generally, non-minimally derivative couplings introduce higher than second-order derivatives in the field equations and extra degrees of freedom, potentially leading to the Boulware-Deser ghost [49]. A particular choice, where the derivative coupling takes the form $G^{\mu\nu}\phi_{,\mu}\phi_{,\nu}$ which is a special case of the most general Horndeski theory [50], avoids this issue because the resulting field equations in Horndeski theory contain no higher than second derivatives [50, 51]. Moreover, the new Higgs inflation model, through the introduction of the non-minimally derivative coupling, satisfies the unitarity bound [52, 53]. The gravitationally enhanced friction resulting from this coupling also slows down the evolution of the scalar field [52, 53, 54, 55]. In this paper, we show that various inflationary models with such non-minimally derivative coupling are well consistent with the latest ACT data.
@@ 2 The model with non-minimally derivative coupling
The action for the scalar field with the kinetic term non-minimally coupled to the Einstein tensor is [51, 52, 53]
@@
where the Einstein tensor $G_{\mu\nu}=R_{\mu\nu}-g_{\mu\nu}R/2$, $R_{\mu\nu}$ is the Ricci tensor, $R$ is the Ricci scalar, $M$ is the coupling constant with the dimension of mass and we take the Planck mass $M_{pl}^{2}=(8\pi G)^{-1}=1$. From the action (2.1), we get the Friedmann equation and the equation of motion for the scalar field $\phi$
@@
where $V_{\phi}=dV/d\phi$. To quantify those slow-roll conditions (2.4) in the high friction limit $H^{2}\gg M^{2}$, we introduce the following slow-roll parameters [53, 55]
@@
where $c_{s}$ and $c_{t}$ are the effective sound speed, $\nu\approx 3/2+4\epsilon-\eta$ and $\mu=3/2+\epsilon$. The amplitude of scalar power spectrum, the scalar spectral tilt and the tensor-to-scalar ratio are [53, 55]
@@
In GR, the slow-roll parameters are defined as
$\epsilon_{v}=\frac{1}{2}\left(\frac{V_{\phi}}{V}\right)^{2},$ (2.14)
$\eta_{v}=\frac{V_{\phi\phi}}{V}.$ (2.15)
The scalar spectral tilt and the tensor-to-scalar ratio are
$n_{s}\approx 1-6\epsilon_{v}+2\eta_{v},$ (2.16)
$r\approx 16\epsilon_{v}.$ (2.17)
@@
$N(\phi_{*})=\int_{\phi_{*}}^{\phi_{e}}\frac{H}{\dot{\phi}}d\phi=\int_{\phi_{*}}^{\phi_{e}}\frac{1}{\sqrt{2\epsilon}}\sqrt{\frac{3H^{2}}{M^{2}}}d\phi,$ (2.18)
where the subscript “ee” means the value of inflaton field at the end of inflation and the subscript “*” means the value at the horizon crossing. The value of $N$ is around $\sim 50-60$.
@@ 3.1 Power law potential
Note that $M^{2}\ll V_{0}$ in the high friction limit, without loss of generality, we take $M^{2}/V_{0}=10^{-4}$ in this paper. The end of inflation is determined by ${\rm max}\{\epsilon(\phi_{e}),|\eta(\phi_{e})|\}=1$. For $0<n<2$, $|\eta|<\epsilon$, inflation ends when $\epsilon(\phi_{e})=1$, so we have
@@
For $n\geq 2$, $|\eta|>\epsilon$, inflation ends when $\eta(\phi_{e})=1$, so
@@
where $\tilde{N}=N+n/2(n+2)$ for $0<n<2$ and $\tilde{N}=N+(n-1)/(n+2)$ for $n\geq 2$. Substituting Eqs. (3.3) and (3.5) into Eqs. (2.12) and (2.13), we obtain the scalar spectral tilt and the tensor-to-scalar ratio,
@@ 3.2 Hilltop model
Since $\phi<\mu$, so $\epsilon(\phi)$ is smaller than $\eta(\phi)$ by a factor $(\phi/\mu)^{p}$ when $p\geq 3$. In this model, we take $p=4$ as a representative case. By using Eqs. (2.12), (2.13) and (2.18), we numerically calculate the observables $n_{s}$ and $r$ with $\mu$ varying and the results are shown in Fig. 1. The solid blue lines represent the results for hilltop model with $p=4$ in the high friction limit, where the left and the right lines correspond to $N=50$ and $N=60$, respectively. In Fig. 1, we also show the results in GR with dashed blue lines. As seen from Fig. 1, the non-minimally derivative coupling increases $n_{s}$ and suppresses $r$, aligning the model with current observational constraints. For $N=50$ the model can be consistent with the P-ACT-LB-BK18 data only at the $2\sigma$ confidence level, while satisfying current observations at the $1\sigma$ confidence level for $N=60$. For $N=50$, we find that $0.18<\mu<2.4$ at the $2\sigma$ confidence level. For $N=60$, we find that $0.165<\mu<0.69$ at the $1\sigma$ confidence level and $\mu>0.123$ at the $2\sigma$ confidence level.
@@
As discussed above, $\epsilon(\phi)$ is smaller than $\eta(\phi)$ by a factor $(\phi/\mu)^{p}$ when $p\geq 3$, so $\epsilon(\phi)<|\eta(\phi)|$, inflation ends when $|\eta(\phi_{e})|=1$. In the limit of $\mu\ll 1$, we get small field inflation, $\phi\ll\mu$,
@@
$\begin{split}{\color[rgb]{0,0,1}\epsilon_{v}}(\phi)&=\frac{p^{2}(\phi/\mu)^{2p-2}}{2\mu^{2}\left[1-(\phi/\mu)^{p}\right]^{2}},\\ {\color[rgb]{0,0,1}\eta_{v}}(\phi)&=-\frac{p(p-1)(\phi/\mu)^{p-2}}{\mu^{2}\left[1-(\phi/\mu)^{p}\right]}=-\frac{2(p-1)}{p}\frac{1-(\phi/\mu)^{p}}{(\phi/\mu)^{p}}{\color[rgb]{0,0,1}\epsilon_{v}}(\phi).\end{split}$ (3.16)
When $p\geq 3$, $\epsilon_{v}(\phi)<|\eta_{v}(\phi)|$, the end of inflation is determined by $|{\color[rgb]{0,0,1}\eta_{v}}(\phi_{e})|=1$ . The number of e-folds before the end of inflation is $N=g(\phi_{e}/\mu)-g(\phi_{*}/\mu)$, where
@@ 3.3 Polynomial $\alpha$-attractors
For the polynomial $\alpha$-attractors [59], the potential is
$V(\phi)=V_{0}\left(1-\frac{\mu^{p}}{\phi^{p}}\right).$ (3.18)
The model is also called D-brane inflation [60]. The slow-roll parameters in the high friction limit are
@@
Without loss of generality, we take $p=3$. By using Eqs. (2.12), (2.13) and (2.18), we numerically calculate the observables $n_{s}$ and $r$ with $\mu$ varying, and the results are shown in Fig. 1 with solid magenta lines. With the help of non-minimally derivative coupling, the polynomial $\alpha$-attractors are consistent with the P-ACT-LB-BK18 data for both $N=50$ and $N=60$. For $N=50$, the $1\sigma$ and $2\sigma$ constraints on $\mu$ by the P-ACT-LB-BK18 data are $0.032<\mu<0.51$ and $0<\mu<3.56$, respectively; while for $N=60$, the $1\sigma$ constraint is $0<\mu<0.57$ and at the $2\sigma$ confidence level, $\mu$ can be arbitrarily large.
In the limit of $\mu\gg 1$, the same attractors (3.7) and (3.8) with $n=1$ are reached because the potential behaves as the linear potential when $\phi\sim\mu$. On the other hand, when $\mu\ll 1$, we have small field inflation. From Eq. (3.19), we see that $\epsilon(\phi)<|\eta(\phi)|$, so $|\eta(\phi_{e})|=1$ and we get
@@
$\begin{split}{\color[rgb]{0,0,1}\epsilon_{v}}(\phi)&=\frac{p^{2}(\mu/\phi)^{2p+2}}{2\mu^{2}\left[1-(\mu/\phi)^{p}\right]^{2}},\\ {\color[rgb]{0,0,1}\eta_{v}}(\phi)&=-\frac{p(p+1)(\mu/\phi)^{p+2}}{\mu^{2}\left[1-(\mu/\phi)^{p}\right]}=-\frac{2(p+1)}{p}\frac{1-(\mu/\phi)^{p}}{(\mu/\phi)^{p}}{\color[rgb]{0,0,1}\epsilon_{v}}(\phi).\end{split}$ (3.25)
@@ 3.4 Exponential $\alpha$-attractors
These models were also called exponential $\alpha$-attractors in [59]. In this section, without loss of generality we consider $a=1$ for simplicity. The slow-roll parameters are
@@
By using Eqs. (2.12), (2.13) and (2.18), we numerically calculate $n_{s}$ and $r$ with varying $\alpha$ and the results are shown in Fig. 1. It is evident that the model is consistent with the P-ACT-LB-BK18 data at the $2\sigma$ confidence level with the help of non-minimally derivative coupling. For $N=50$, the $1\sigma$ and $2\sigma$ constraints on $\alpha$ are $1.2\times 10^{-3}<\alpha<0.011$ and $2.5\times 10^{-4}<\alpha<0.468$, respectively. For $N=60$, the $1\sigma$ and $2\sigma$ constraints are $1.38\times 10^{-4}<\alpha<0.022$ and $\alpha>0$, respectively.
@@
In the small $\alpha$ limit, $\alpha\ll 1$, the end of inflation is determined by $|\eta(\phi_{e})|=1$ because $\epsilon(\phi)<|\eta(\phi)|$, so we have
@@
$\begin{split}{\color[rgb]{0,0,1}\epsilon_{v}}(\phi)&=\frac{1}{3\alpha\left[\exp\left(\sqrt{\frac{2}{3\alpha}}\phi\right)-1\right]^{2}},\\ {\color[rgb]{0,0,1}\eta_{v}}(\phi)&=\frac{2}{3\alpha\left[1-\exp\left(\sqrt{\frac{2}{3\alpha}}\phi\right)\right]}.\end{split}$ (3.35)
@@ 4 Conclusions
For the hilltop model, the scale $\mu$ can be smaller than the Planck energy, allowing for small-field inflation. In the case $p=4$, the scenario with $N=50$ lies within the $2\sigma$ $n_{s}-r$ contour from the P-ACT-LB-BK18 data, while the case with $N=60$ falls within both the $1\sigma$ and $2\sigma$ $n_{s}-r$ contours. For $N=50$, we constrain the scale as $0.18<\mu<2.4$ at the $2\sigma$ confidence level. For $N=60$, the bounds are $0.165<\mu<0.69$ at $1\sigma$, with a lower limit $\mu>0.123$ at $2\sigma$. Additionally, we derive analytical expressions for attractor behaviors in the limits of large and small $\mu$, for both the non-minimal derivative coupling scenario and GR.
For the polynomial $\alpha$-attractors with $p=3$, both the $N=50$ and $N=60$ cases are consistent with the P-ACT-LB-BK18 data. For $N=50$, the $1\sigma$ and $2\sigma$ constraints on $\mu$ are $0.032<\mu<0.51$ and $0<\mu<3.56$, respectively. For $N=60$, the $1\sigma$ constraint is $0<\mu<0.57$, while at the $2\sigma$ level, $\mu$ is unconstrained and can be arbitrarily large. The numerical results confirm the analytical attractor expressions for both the non-minimally derivative coupling scenario and GR in the limits of large and small $\mu$.